OpenStudy (anonymous):

find the slope of (-7 and 5)

OpenStudy (anonymous):

is there any other point, or are they on the y or x axis?

OpenStudy (anonymous):

they are on the y and x axis

OpenStudy (anonymous):

but thats just one point, the slope is the incline

OpenStudy (anonymous):

oh ok let me look again

OpenStudy (anonymous):

sorry (-3,1) and (5,-7)

OpenStudy (anonymous):

its fine; and the formula for the slope it y2 - y1 m = ------ x2 - x1 -7 - 1 m = ------ 5 - (-3) -8 m = ------ 8 m = -1

OpenStudy (anonymous):

m = the slope by the way; that means the slope is -1

OpenStudy (anonymous):

got it, thank you

OpenStudy (anonymous):

np (: